A solid whose mass is 13.5 g, is placed in a cylinder with water. The initial volume of water is 60 mL. The solid displaced the water and it reached the level of 67.5 mL. Calculate the density of the solid in g/cm3.

Group of answer choices





1
Expert's answer
2021-10-18T04:51:30-0400

Volume of solid"= 67.5-60=7.5ml"


Density "=\frac{13.5}{7.5}=1.8g/cm^3"